A relaxed spring of length 0.17 m stands vertically on the floor; its stiffness is 1460 N/m. You release a block of mass 0.6 kg from rest, with the bottom of the block 0.4 m above the floor and straight above the spring. How long is the spring when the block comes momentarily to rest on the compressed spring?
Answer #1

Free length, L = 0.17 m

Using law of conservation of energy,

0.5 kx^2 = 0.5 mgh

0.5 x 1460 x x^2 = 0.6 x 9.8 x (0.23 + x)

730 x^2 - 5.88 x - 1.35 = 0

X = 0.047 m

Spring length = 0.17 - 0.047 = 0.123 m

Comment in case any doubt please rate my answer ...